A sprinter reaches his maximum speed in 2.5 seconds from rest with constant acceleration.He mantains
this speed and finishes 100 meters in over all time of 10.40 seconds. find his maximum speed.
1
Expert's answer
2012-04-05T09:07:54-0400
V=at
x=at^2/2 <=>100=a*54,08 =>
a=100/54.8
Vmax=a*tmax=100*2.5/54.8=4.57